OpenStudy (littleheron):

Need help fast! Will Medal! Calculate the slope of the line given the points (-2, 6) and (0, 4). -1 1 0 none of the above

Mehek (mehek14):

\(\dfrac{4-6}{0-(-2)}=?\)

OpenStudy (littleheron):

.-.

Mehek (mehek14):

4-6 = ?

OpenStudy (littleheron):

-2

Mehek (mehek14):

0-(-2)=0+2 = ?

OpenStudy (littleheron):

?

Mehek (mehek14):

0+2= ?

OpenStudy (littleheron):

2

Mehek (mehek14):

-2/2 = ?

OpenStudy (littleheron):

-1

Mehek (mehek14):

so slope = -1

OpenStudy (littleheron):

o-o that was suprisingly easy

Mehek (mehek14):

yes :P just use the formula \(\dfrac{y_2-y_1}{x_2-x_1}\)